OpenStudy (anonymous):

f(x)=((x-2)/(3)) g(x)= x^2 Find (f o g)(x).

OpenStudy (john_es):

\[(f\circ g)(x)=f(g(x))=\frac{g(x)-2}{3}=\frac{x^2-2}{3}\]

OpenStudy (anonymous):

Thank you! I just wanted to make sure I had done it right. How do I find the domain for that function?

OpenStudy (john_es):

It is a polinomy, so its domain is "all reals"

OpenStudy (john_es):

If the function was, \[\frac{2}{x^2-4}\]Then as it is a rational function, you should equate the denominator to 0, \[x^2-4=0\Rightarrow x=\pm2\] As a fraction cannot have a 0 denominator, then its domain is all reals except -2 and +2.